• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name slifer

    Log entry time 05:16:55 on October 15,1998

    Entry number 2923

    keyword=dipoles

    The regulator for the dipoles apparently is not able to operate properly at the
    present momentum setting.(.447 Gev) so the regulators
    were turned off and the currents set by hand.
    Electron:146.75A
    Hadron :146.20A
    The Hadron quickly deviates out of specs so we retained the
    prescaler that the previous shift had set.
    NMR measurements at 04:00 and 05:08 verified that the electron dipole
    was at the correct field with a deviation < 1*10^-3